# thanks: Current address: Laboratoire de Physique de l’École normale supérieure, Université PSL, CNRS, Sorbonne Université and Université Paris Cité, 75005 Paris, France

### v1

@@ III.5 Characterizing the wetting transitions
Using the simulation data, we now characterize the type of wetting transition, which is crucial for our understanding of this phenomenon. For equilibrium wetting transitions on a solid-liquid interface, both first-order and continuous transitions have been predicted and observed, depending on the types of interactions between the components and the system temperature cahn1977critical ; nakanishi1982multicriticality ; wang2001line ; murata2010surface . For active systems, an appropriate order parameter has to be defined to track such transition properties. Sepúlveda and Soto have studied the wall aggregation of active particles on inert walls as a function of particle reorientation rate in the bulk sepulveda2017wetting ; SepulvedaSoto_PRE2018_activewetting , and the instantaneous fraction of empty sites on the wall has been used as the order parameter. Similarly, we use the vacancy fraction $\langle\ell\rangle$ as the order parameter here, which has been calculated by averaging the vacancy fraction over time and for multiple realizations for a given set of parameters. As was done in previous studies, we compute $\langle\ell\rangle$ as a function of $Pe$, which is a measure of active particle persistence. In Fig. 5, we plot $\langle\ell\rangle$ vs $Pe^{-1}$ for the range of $D_{r}^{w}$ where partial wetted states are also observed. For a small Pe (large Pe-1) the system displays a complete dry state as $\langle\ell\rangle\rightarrow 1$ and for larger Pe (smaller Pe-1), $\langle\ell\rangle$ is between 0 and 1. The transition is rather discontinuous for relatively small $D_{r}^{w}$, as the change in $\langle\ell\rangle$ is rather abrupt. However, when $D_{r}^{w}$ is increased, we observe signs of a continuous transition from $\langle\ell\rangle<1$ to $\langle\ell\rangle\simeq 1$. In this range, $\langle\ell\rangle$ even indicates a power-law-like behavior as a function of $Pe$. For $D_{r}^{w}\simeq 375$, $\ell$ vs $Pe^{-1}$ approximately fits to a power-law, $\ell\sim Pe^{-\alpha}$, where $\alpha\simeq 1$ (see Fig. 5). Interestingly, the exponent is consistent with previous observations on wetting transitions in active systems sepulveda2017wetting ; SepulvedaSoto_PRE2018_activewetting . However, the transition that we observe is from a partial wetted to a complete dry state (drying transition), unlike shown in reference sepulveda2017wetting . In those studies, instead of changing the wall properties, the wetting transitions were observed by changing the particle reorientation rate. Therefore, this observation is indicating the existence of a universal behavior in wetting transitions achieved through various mechanisms.
@@ IV Summary and discussions
The study by Rogers et al. Rogers_PRL_2012 on the rotational diffusion of colloids near surfaces offers valuable experimental inspiration. A potential experimental validation of our findings could involve bacteria confined within rigid walls treated with chemical repellents. Adjusting the repellent strength would influence bacterial movement away from the wall, providing a controlled way to study wall-particle interactions.
@@ Appendix B Steady-state density distribution and Vacancy fraction
To characterize the steady-state density distribution, we divide the system into $b$ square bins and measure the time-averaged number of particles in each bin, $\langle n_{b}\rangle$. The two dimensional density distribution is then given by $\rho(x,y)=\langle n_{b}\rangle\pi\sigma^{2}/(4A_{b})$, where $A_{b}$ denotes the area of the bin, chosen as $A_{b}=2.0r_{c}$. We plot this in Fig. 1(b) for a typical steady-state configuration. For small $D_{r}^{w}$, we observe high local density near both the walls as shown in Fig. 1(b)(i), corresponding to a symmetric complete wetting state. Increasing $D_{r}^{w}$, we observe high local density only on one wall, as shown in Fig. 1(b)(ii). This transition occurs around $D_{r}^{w}=135$, corresponding to an asymmetric wetting state. Further increase of $D_{r}^{w}$, we observe droplet formation as shown in Fig. 1(b)(iii) for $D_{r}^{w}=375$, corresponding to a partial wetting state.
To quantify the different morphological states, we calculate the vacancy fraction $\ell$. We bin the system along the $x$-axis and calculate the number of unoccupied bins, normalized by the system length along the $x$ axis. From this, we calculate the fraction of vacant length along the wall and take the mean over both walls. This quantity is defined as the vacancy fraction, $\ell=(\ell_{1}+\ell_{2})/2$, where $\ell_{1}$ and $\ell_{2}$ are the vacancy fraction for the lower and upper wall, respectively, as shown in Fig.A1. In Fig.A2, we show typical examples for the four different wetting states.
Figure A1: Schematic for calculating vacancy fraction $\ell=(\ell_{1}+\ell_{2})/2$.
Figure A2: Example of vacancy fraction for (a) symmetric complete wetting state ($D_{r}^{w}=30$), (b) asymmetric complete wetting ($D_{r}^{w}=135$), (c) partial wetting ($D_{r}^{w}=375$) and (d) complete drying state ($D_{r}^{w}=960$) with $N=38400$, $Pe=240$.
@@ Appendix D Bi-stability statistics
We quantify the number of independent runs showing the different wetting states, where the transition regions between the states show indications of bistability, as shown in Fig.A3. We note that for some parameters, we have observed late-time transitions from one morphological state to another. Thus, the characterization of the morphological states is based on the transitions observed until the end of the simulation.
Figure A3: Fraction of runs showing the different wetting states, with transition regions showing bistability, characterised here for a system having $N=8008$ particles for $Pe=150$, for $80$ independent runs.
@@ Appendix E Estimation of contact angle and its variation with activity
To quantify the effect of activity on droplet morphology, we calculate the contact angle $\theta$ as a function of activity $Pe$. Our analysis confirms that $\theta$ increases with activity, as shown in Fig.A4, consistent with the increase in interfacial curvature.
Figure A4: Variation in contact angle of droplet with activity for $D_{r}^{w}=540$, corresponding to Fig. 3(a).
@@ Appendix G Quantification of bubbles
Figure A5: Typical snapshot of observed bubble formation, shown here for $D_{r}^{w}=120$, $Pe=150$.
Figure A6: (a) Local density distribution $P(\varphi)$ vs. $\varphi$ for increasing $D_{r}^{w}$ plotted in log-linear scale. $P(\varphi)$ follows the normalization $\sum_{i}p(\varphi_{i})\Delta\varphi_{i}=1$. (b) Zoomed in view highlighting the lower tail of $P(\varphi)$, plotted on a linear-linear scale.
Appendix H Supplementary movies
  • MOVIE1: Typical cluster morphology observed for symmetric complete wetting (SCW) state, observed at $Pe=240$ and $D_{r}^{w}$= 30.
  • MOVIE2: Typical cluster morphology observed for asymmetric complete wetting (ACW) state, observed at $Pe=240$ and $D_{r}^{w}$= 135.
  • MOVIE3: Typical cluster morphology observed for partial wetting (PW) state, observed at $Pe=240$ and $D_{r}^{w}$= 375.

### v2

@@ III.5 Characterizing the wetting transitions
Using the simulation data, we now characterize the type of wetting transition, which is crucial for our understanding of this phenomenon. For equilibrium wetting transitions on a solid-liquid interface, both first-order and continuous transitions have been predicted and observed, depending on the types of interactions between the components and the system temperature cahn1977critical ; nakanishi1982multicriticality ; wang2001line ; murata2010surface . For active systems, an appropriate order parameter has to be defined to track such transition properties. Sepúlveda and Soto have studied the wall aggregation of active particles on inert walls as a function of particle reorientation rate in the bulk sepulveda2017wetting ; SepulvedaSoto_PRE2018_activewetting , and the instantaneous fraction of empty sites on the wall has been used as the order parameter. Similarly, we use the vacancy fraction $\langle\ell\rangle$ as the order parameter here, which has been calculated by averaging the vacancy fraction over time and for multiple realizations for a given set of parameters. As was done in previous studies, we compute $\langle\ell\rangle$ as a function of $Pe$, which is a measure of active particle persistence. In Fig. 5, we plot $\langle\ell\rangle$ vs $Pe^{-1}$ for the range of $D_{r}^{w}/D_{r}$ where partial wetted states are also observed. For a small Pe (large Pe-1) the system displays a complete dry state as $\langle\ell\rangle\rightarrow 1$ and for larger Pe (smaller Pe-1), $\langle\ell\rangle$ is between 0 and 1. The transition is rather discontinuous for relatively small $D_{r}^{w}/D_{r}$, as the change in $\langle\ell\rangle$ is rather abrupt. However, when $D_{r}^{w}$ is increased, we observe signs of a continuous transition from $\langle\ell\rangle<1$ to $\langle\ell\rangle\simeq 1$. In this range, $\langle\ell\rangle$ even indicates a power-law-like behavior as a function of $Pe$. For $D_{r}^{w}/D_{r}\simeq 125$, $\ell$ vs $Pe^{-1}$ approximately fits to a power-law, $\ell\sim Pe^{-\alpha}$, where $\alpha\simeq 1$ (see Fig. 5). Interestingly, the exponent is consistent with previous observations on wetting transitions in active systems sepulveda2017wetting ; SepulvedaSoto_PRE2018_activewetting . However, the transition that we observe is from a partial wetted to a complete dry state (drying transition), unlike shown in reference sepulveda2017wetting . In those studies, instead of changing the wall properties, the wetting transitions were observed by changing the particle reorientation rate. Therefore, this observation is indicating the existence of a universal behavior in wetting transitions achieved through various mechanisms.
@@ IV Summary and discussions
The study by Rogers et al. Rogers_PRL_2012 on the rotational diffusion of colloids near surfaces offers valuable experimental inspiration. Recently developed self-propelled robots can be used to validate our numerical findings paramanick2024programming . Another potential experimental application of our findings could involve bacteria confined within rigid walls treated with chemical repellents. Adjusting the repellent strength would influence bacterial movement away from the wall, providing a controlled way to study wall-particle interactions.
Our study on the wetting behavior of active Brownian particles can be extended in many different directions. For example, anisotropic or rod-like self-propelled objects are expected to have a different wetting-like behavior which can be systematically explored wensink2008aggregation ; elgeti2009self ; khan2022effect . Further, the wall aggregation of deformable particles loewe2020solid ; ohta2009deformable , and active ring polymers mousavi2019active ; theeyancheri2024dynamic can also be systematically explored. This study can be repeated for different types of walls with varying curvature, roughness, geometry, and in the presence of external fields Bechinger2016 ; pincce2016disorder ; shaik2023confined . Our study is conducted for a fixed area fraction (or a fixed number of particles), and the morphological changes are induced due to the change in wall interaction. One pertinent question is how this wetting behavior will be influenced if the number of particles is not constant, in a grand canonical arrangement neta2021wetting .
Supplementary Material
For list of parameters, supplementary figures and associated text, movie captions, refer to supplementary material.
@@
Author declarations
Conflict of interest: The authors have no conflicts to disclose.
Author contributions- SD and RC: Conceptualization (equal); developing the numerical model (equal); running the simulations (equal); Analyzing the data (equal); Writing the manuscript (equal).
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Supplemental Material for “Active wetting transitions induced by rotational noise at solid interfaces”
@@ Appendix B Steady-state density distribution and Vacancy fraction
To characterize the steady-state density distribution, we divide the system into $b$ square bins and measure the time-averaged number of particles in each bin, $\langle n_{b}\rangle$. The two dimensional density distribution is then given by $\rho(x,y)=\langle n_{b}\rangle\pi\sigma^{2}/(4A_{b})$, where $A_{b}$ denotes the area of the bin, chosen as $A_{b}=2.0r_{c}$. We plot this in Fig. 1(b) for a typical steady-state configuration. For small $D_{r}^{w}/D_{r}$, we observe high local density near both the walls as shown in Fig. 1(b)(i) corresponding to a symmetric complete wetting state. Increasing $D_{r}^{w}/D_{r}$, we observe high local density only on one wall, as shown in Fig. 1(b)(ii). This transition occurs around $D_{r}^{w}/D_{r}=45$, corresponding to an asymmetric wetting state. Further increase of $D_{r}^{w}/D_{r}$, we observe droplet formation as shown in Fig. 1(b)(iii) for $D_{r}^{w}/D_{r}$ = 125, corresponding to a partial wetting state.
To quantify the different morphological states, we calculate the vacancy fraction $\ell$. We bin the system along the $x$-axis and calculate the number of unoccupied bins, normalized by the system length along the $x$ axis. From this, we calculate the fraction of vacant length along the wall and take the mean over both walls. This quantity is defined as the vacancy fraction, $\ell=(\ell_{1}+\ell_{2})/2$, where $\ell_{1}$ and $\ell_{2}$ are the vacancy fraction for the lower and upper wall, respectively, as shown in Fig. S1. In Fig. S2, we show typical examples for the four different wetting states.
Figure S1: Schematic for calculating vacancy fraction $\ell=(\ell_{1}+\ell_{2})/2$.
Figure S2: Example of vacancy fraction for (a) symmetric complete wetting state ($D_{r}^{w}/D_{r}=10$), (b) asymmetric complete wetting ($D_{r}^{w}/D_{r}=45$), (c) partial wetting ($D_{r}^{w}/D_{r}=125$) and (d) complete drying state ($D_{r}^{w}/D_{r}=320$) with $N=38400$, $Pe=240$.
@@ Appendix D Bi-stability statistics
We quantify the number of independent runs showing the different wetting states, where the transition regions between the states show indications of bistability, as shown in Fig. S3. We note that for some parameters, we have observed late-time transitions from one morphological state to another. Thus, the characterization of the morphological states is based on the transitions observed until the end of the simulation.
Figure S3: Fraction of runs showing the different wetting states, with transition regions showing bistability, characterized here for a system having $N=8008$ particles for $Pe=150$, for $80$ independent runs.
@@ Appendix E Estimation of contact angle and its variation with activity
To quantify the effect of activity on droplet morphology, we calculate the contact angle $\theta$ as a function of activity $Pe$. Our analysis confirms that $\theta$ increases with activity, as shown in Fig. S4, consistent with the increase in interfacial curvature.
Figure S4: Variation in contact angle of droplet with activity for $D_{r}^{w}/D_{r}=180$, corresponding to Fig. 3(a).
@@ Appendix G Quantification of bubbles
Figure S5: Typical snapshot of observed bubble formation, shown here for $D_{r}^{w}/D_{r}=40$, $Pe=150$.
Figure S6: (a) Local density distribution $P(\varphi)$ vs. $\varphi$ for increasing $D_{r}^{w}/D_{r}$ plotted in log-linear scale. $P(\varphi)$ follows the normalization $\sum_{i}p(\varphi_{i})\Delta\varphi_{i}=1$. (b) Zoomed in view highlighting the lower tail of $P(\varphi)$, plotted on a linear-linear scale.
Appendix H Noninteracting particles
We have conducted simulation of active Brownian particles in the absence of interparticle interactions and compared with the results obtained for interacting ABPs. Fig. S7 shows the density profiles in the absence of particle interaction ${\bf{F}}_{i}=0$, for the parameters in which we observe (a) symmetric, and (b) asymmetric complete wetting for interacting particles. The particle aggregation still occurs even in the absence of interparticle interactions. However, we do not observe an asymmetric complete wetting state in this case, as the liquid-vapour interface is absent.
Figure S7: Comparison of density profiles with and without particle interaction, for (a) symmetric wetting case $D_{r}^{w}/D_{r}=10$, $Pe=240$, (b) asymmetric wetting case $D_{r}^{w}/D_{r}=45$, $Pe=240$.
Appendix I Comparison of the area fraction
We compare the area fractions observed in the dense and vapor states in both the symmetric and asymmetric complete wetting states, and contrast these with the case of active Brownian particles (ABPs) undergoing motility-induced phase separation (MIPS), as shown in Fig. S8. To distinguish between the liquid and vapor phases, we analyze the mean interface height, $\langle h\rangle$, which quantifies the location of the liquid-vapor interface. We then compute the averaged density profiles separately for the dense and vapor phases and compare them with the corresponding profiles for ABPs undergoing MIPS, as reported in reference Siebert2018 .
Figure S8: Comparison of area fractions of the liquid and vapor states in (i) the case of active Brownian particles in periodic boundary conditions (data extracted from Siebert et al. Siebert2018 ., (ii) SCW state ($D_{r}^{w}/D_{r}=10$, $Pe=150$ and $Pe=240$), and (iii) ACW state ($D_{r}^{w}/D_{r}=45$, $Pe=150$ and $Pe=240$)
Appendix J Supplementary movies
  • MOVIE1: Typical cluster morphology observed for symmetric complete wetting (SCW) state, observed at $Pe=240$ and $D_{r}^{w}/D_{r}$ = 10.
  • MOVIE2: Typical cluster morphology observed for asymmetric complete wetting (ACW) state, observed at $Pe=240$ and $D_{r}^{w}/D_{r}$ = 45.
  • MOVIE3: Typical cluster morphology observed for partial wetting (PW) state, observed at $Pe=240$ and $D_{r}^{w}/D_{r}$ = 125.